Given this matrix, write the Characteristic Equation, and the Eigenvalues with their proper eigenvectors

A = [ 17 21 2 4
Answer #1

A = 72] Page 1 To find characteristic squation, | A-20]=0 113:3-46938 al 7-> 2 2 4- (7-2)(4-2) – 4 =0 X-112+28 - 4 =0 X2_11xNext, find the eigen vectors, 2 Page 2 at x = 8 (A-XI) = (A - 81) (51)-(3) · 1)?:) row operations reduce the above matrix byAt x=3 (A-NE) - (A-32) Page 3 -:-) 4 2 i 2 How operations to obtain reduced matrix Perform, (R2 + R2 - RO) 4 2 0 O (R. - ) co